what is the voulume if 34.0 g of ammonia (NH3) at stp?
1
Expert's answer
2018-02-05T07:23:09-0500
V(NH3) = n × Vm, where n is the amount of substance (mole), Vm is the molar volume (litre/mole). For any gas 1 mole will occupy 22.4 litre.

n = m/M, where m is the mass of substance (mole), M is the molar mass (g/mole)

V(NH3) = m × Vm / M

V(NH3) = 34 × 22.4 / 17 = 44.8 litre